Can you please help me to understand how Re=1-(1-R)(1-R) becomes Re=2R-R^2. I multiplied out the brackets and was left with 1-1-R-R+R^2, which gave me -2R+R^2
Answer by josmiceli(19441)

Re=1-(1-R)(1-R)
Re = 1 - ( 1 - 2R + R^2 )
Re = 1 - 1 + 2R - R^2
Re = 2R - R^2